# Accretion of a Vlasov gas by a Kerr black hole

### v1

@@ Abstract
We investigate the accretion of a collisionless, relativistic kinetic gas by a rotating Kerr black hole, assuming that at infinity the state of the gas is described by a distribution function depending only on the energy of the particles. Neglecting the self-gravity of the gas, we show that relevant physical observables, including the particle current density and the accretion rates associated with the mass, the energy, and the angular momentum, can be expressed in the form of closed integrals that can be evaluated numerically or approximated analytically in the slow-rotation limit. The accretion rates are computed in this manner for both monoenergetic particles and the Maxwell-Jüttner distribution and compared with the corresponding results in the non-rotating case. Whereas it is shown that the angular momentum accretion rate vanishes exactly, it is found that the rotation of the black hole has a small but non-vanishing effect on the mass and the energy accretion rates, which is remarkably well described by an analytic calculation in the slow-rotation approximation to quadratic order in the rotation parameter. The effects of rotation on the morphology of the accretion flow are also analyzed.
@@ I Introduction
At the same time, non-spherical flows in the Schwarzschild spacetime were investigated within the same framework in [18, 19, 20] (axially symmetric disks), or [21, 22, 23, 24] (accretion onto moving black holes). A Bondi-type model confined to the equatorial plane of the Kerr spacetime was derived in [25] and [26]. Analytic kinetic models describing the relaxation (due to phase-space mixing) of macroscopic variables to a stationary state around a Kerr black hole have been described in [27, 28]. For self-gravitating shells or disks made of Vlasov matter around black holes, see [29, 30, 31, 32].
@@
We show that the particle number and energy accretion rates decrease with the increasing black hole spin parameter. This stays in agreement with previous results, in particular the model of accretion onto Reissner-Nordström black holes [11, 12] and the planar model discussed in [25]. This is also a general feature present in hydrodynamical models, e.g., in [5]. In contrast to the results of [25], both accretion rates decrease with increasing asymptotic temperature of the gas. This fact emphasizes the difference between planar and non-planar models, mostly due to their asymptotic properties. Perhaps more surprisingly, we show that the total angular momentum accretion rate vanishes. This stays in agreement with the hydrodynamical model of [1], but not with the properties of the planar model [25].
Our solution is exact in the sense that the relevant observable quantities (such as the components of the particle current density or the components of the energy-momentum tensor) can be expressed in the form of closed integrals. These integrals are then evaluated numerically. As an example, we compute the components of the particle current density, illustrating the morphology of the flow in the vicinity of the black hole.
@@ II Preliminaries
$\hat{L}_{z}:=L_{z}-aE,\qquad\beta:=\frac{\hat{L}_{z}}{L}.$ (8)
@@ III.2 New parametrization
    In the Schwarzschild limit $\alpha=0$, $Q_{c}=L_{c}(E)$, and $Q_{\mathrm{max}}=L_{\mathrm{max}}(r,E)$ are independent of $\alpha$ and $\sin\chi$ and reduce to the corresponding expressions in [9].
@@ III.5 Asymptotic conditions
$F_{\infty}=Am\delta(E-E_{0})=A\delta(\varepsilon-\varepsilon_{0}).$ (61)
In this case, the gas comprises of same-energy particles. Our second choice would be to set
@@ IV Accretion rates
The energy and angular momentum accretion rates can be obtained in the same fashion by inserting the factors $p_{t}=E$ and $p_{\varphi}=Q\cos\chi$ under the integral (75), which yields
@@
$\mathcal{\dot{J}}=-\frac{2\pi}{3}\int\limits_{m}^{\infty}dEF_{\infty}(E)\int\limits_{0}^{\pi}d\vartheta\sin\vartheta\int\limits_{0}^{2\pi}d\chi\cos\chi(Q_{c})^{3}=0,$ (78)
@@
$\frac{\dot{\mathcal{N}}}{n_{\infty}}=\frac{\dot{\mathcal{E}}}{\varepsilon_{\infty}}=-\frac{M^{2}}{4\varepsilon_{0}\sqrt{\varepsilon_{0}^{2}-1}}\int\limits_{0}^{\pi}d\vartheta\sin\vartheta\int\limits_{0}^{2\pi}d\chi q_{c}(\alpha,\vartheta,\sin\chi,\varepsilon_{0})^{2},$ (80)
@@
$\lim\limits_{z\to\infty}\frac{1}{\sqrt{2\pi z}}\frac{\dot{\mathcal{N}}}{n_{\infty}}=\lim\limits_{z\to\infty}\frac{1}{\sqrt{2\pi z}}\frac{\dot{\mathcal{E}}}{\varepsilon_{\infty}}=-\frac{M^{2}}{4\pi}\int\limits_{0}^{\pi}d\vartheta\sin\vartheta\int\limits_{0}^{2\pi}d\chi q_{c}(\alpha,\vartheta,\sin\chi,1)^{2}.$ (82)
In the high-temperature limit one uses the substitution $\varepsilon\mapsto\eta:=z\varepsilon$ and $K_{\nu}(z)\sim\Gamma(\nu)(2/z)^{\nu}$ to obtain
$\lim\limits_{z\to 0}\frac{\dot{\mathcal{N}}}{n_{\infty}}=\lim\limits_{z\to 0}\frac{\dot{\mathcal{E}}}{\varepsilon_{\infty}}=-\frac{M^{2}}{4}\int\limits_{0}^{\pi}d\vartheta\sin\vartheta\int\limits_{0}^{2\pi}d\chi\zeta_{c}(\alpha,\vartheta,\sin\chi)^{2},$ (83)
@@ V Slow-rotation approximation
Since $Q_{c}=Q_{c}(a,\vartheta,\chi,E)$ is only given in parametric form (see Eq. (129)), the integrals with respect to $\chi$ present in Eqs. (45) cannot be computed analytically and one has to resort to numerical methods. Although we shall compute these integrals numerically in the upcoming section, here we circumvent this problem by considering a slowly rotating Kerr black hole and performing an expansion in the rotation parameter, up to quadratic order. The slow-rotation approximation is an important astrophysically motivated case [47], and it provides an effective strategy to overcome difficulties arising in the Kerr background spacetime due to the frame-dragging effect [48]. In principle, the analytic formulas presented in this section can be extended to higher order of the black hole spin parameter, to derive relations valid for more rapidly rotating black holes. However, as we will see, the quadratic approximation already leads to an accurate description of the accretion process.
Up to the second order in $\alpha$, the critical value $Q_{c}$ can be expressed explicitly in terms of $E$, $\chi$, and $\vartheta$ in the following way. First, we expand the critical radius
$x_{c}=x_{c}^{(0)}+\alpha x_{c}^{(1)}+\alpha^{2}x_{c}^{(2)}+\mathcal{O}(\alpha^{3}),$ (85)
@@
which is a function of energy only. To find the correction terms $x_{c}^{(1)}$ and $x_{c}^{(2)}$, we substitute the ansatz (85) into the expression for $E_{c}$ given in Eq. (129) and then solve the equation $E_{c}=E$ order by order in $\alpha$. After some manipulations, this yields
@@
Next, we substitute this function into the expression for $Q_{c}$ given in Eq. (129) and expand the result in powers of $\alpha$. In terms of the dimensionless quantities $q_{c}=Q_{c}/(Mm)$ this yields, up to the second order,
$q_{c}=q_{c}^{(0)}+\alpha q_{c}^{(1)}+\alpha^{2}q_{c}^{(2)}+\mathcal{O}(\alpha^{3}),$ (88)
@@
$q_{c}^{(2)}$ $=$ $\frac{q_{c}^{(0)}}{2x_{c}^{(0)}}\left[-\frac{\cos^{2}\vartheta+x_{c}^{(0)}\sin^{2}\vartheta}{x_{c}^{(0)}}+\frac{\left(x_{c}^{(0)}-5\right)\sin^{2}\vartheta\sin^{2}\chi}{x_{c}^{(0)}-6}\right].$ (89c)
Note that the zeroth order term again agrees with the value known from the Schwarzschild case, see Eq. (22). Together with Eq. (86), Eqs. (88) and (89) provide the desired expansion for $Q_{c}$ as a function of $E$, $\chi$, and $\vartheta$.
@@
$\int_{0}^{\pi}d\vartheta\sin\vartheta\int_{0}^{2\pi}d\chi q_{c}^{2}=4\pi\frac{\left(x_{c}^{(0)}\right)^{2}}{x_{c}^{(0)}-3}\left[1-\frac{2\alpha^{2}}{\left(x_{c}^{(0)}\right)^{2}\left(6-x_{c}^{(0)}\right)}+\mathcal{O}(\alpha^{3})\right],$ (90)
@@
On the other hand, one can numerically perform the following integrations over the energy and compute the accretion rates for generic values of $z$ up to the second order in the rotation parameter $\alpha$. Using Eq. (90), one obtains
$\frac{\dot{\mathcal{N}}}{n_{\infty}}$ $=$ $-\frac{\pi zM^{2}}{K_{2}(z)}\int_{1}^{\infty}\frac{\left(x_{c}^{(0)}\right)^{2}}{x_{c}^{(0)}-3}\left[1-\frac{2\alpha^{2}}{\left(x_{c}^{(0)}\right)^{2}\left(6-x_{c}^{(0)}\right)}\right]e^{-z\varepsilon}d\varepsilon+\mathcal{O}(\alpha^{3}),$ (92a)
$\frac{\dot{\mathcal{E}}}{\varepsilon_{\infty}}$ $=$ $-\frac{\pi z^{2}M^{2}}{zK_{1}(z)+3K_{2}(z)}\int_{1}^{\infty}\frac{\left(x_{c}^{(0)}\right)^{2}}{x_{c}^{(0)}-3}\left[1-\frac{2\alpha^{2}}{\left(x_{c}^{(0)}\right)^{2}\left(6-x_{c}^{(0)}\right)}\right]\varepsilon e^{-z\varepsilon}d\varepsilon+\mathcal{O}(\alpha^{3}).$ (92b)
Therefore, within the slow-rotation approximation, the original triple integrals appearing in Eqs. (81) can be replaced with simpler formulas (92). Remarkably, for all numerical examples dicussed in Sec. VI, the difference between the results of the slow-rotation approximation and exact numerical values computed according Eqs. (81) stays below $4\%$, even for rapidly rotating black holes with rotation parameters up to $\alpha\leq 0.99$.
@@
$\lim\limits_{z\rightarrow\infty}\frac{1}{\sqrt{2\pi z}}\frac{J_{t}^{\mathrm{(tot)}}}{n_{\infty}}$ $=$ $-\frac{1}{4\pi x\sqrt{2x}}\left[\frac{16x}{x-2}-\left(1+\frac{4(x+2)\cos^{2}\vartheta}{(x-2)^{2}}\right)\alpha^{2}\right]$ (94a)
$-\frac{1}{4\pi x}\sqrt{\frac{2}{x}}\left[\frac{4(x-4)x}{x-2}+\left(1+\frac{2(x+2)\cos^{2}\vartheta}{(x-2)^{2}}\right)\alpha^{2}\right]1_{x>x_{c},q_{\max}>q_{c}}+\mathcal{O}(\alpha^{3}),$
@@
$\lim\limits_{z\rightarrow\infty}\frac{1}{\sqrt{2\pi z}}\frac{J_{\varphi}^{\mathrm{(tot)}}}{n_{\infty}}$ $=$ $-\frac{2\sqrt{2x}\sin^{2}\vartheta}{\pi(x-2)^{2}}\alpha(1-1_{x>x_{c},q_{\max}>q_{c}})+\mathcal{O}(\alpha^{3}),$ (94c)
where $x_{c}$, $q_{c}$, and $q_{\mathrm{max}}$ in the characteristic function $1_{x>x_{c},q_{\max}>q_{c}}$ should be evaluated at $E=m$.
@@
$\lim\limits_{z\rightarrow\infty}\frac{1}{\sqrt{2\pi z}}\frac{n(x,\vartheta)}{n_{\infty}}=\frac{\sqrt{2c_{0}}}{(x-2)^{2}x^{2}\pi}\left(1+\frac{c_{1}+c_{2}\sin^{2}\vartheta+c_{3}1_{x>x_{c},q_{\max}>q_{c}}}{8c_{0}x}\alpha^{2}\right)+\mathcal{O}(\alpha^{3}),$ (95)
where the parameters $c_{i}$ are defined as
$c_{0}$ $=$ $x(x-2)\left[32(x-1)+4x^{2}(x-2)+x^{4}(x-4)1_{x>x_{c},q_{\max}>q_{c}}\right],$ (96a)
$c_{1}$ $=$ $-2(x^{6}-2x^{5}+24x^{4}-40x^{3}+176x^{2}-256x+128),$ (96b)
$c_{2}$ $=$ $8(x^{5}+8x^{3}+16x^{2}-48x+32),$ (96c)
$c_{3}$ $=$ $x^{4}(x^{3}-4x^{2}+4x+48\cos^{2}\vartheta).$ (96d)
We would like to stress that other observables and physical quantities can be computed within the slow-rotation approximation in an analogous way; the limited examples provided in this work are intended to demonstrate this technique.

### v3

@@ Abstract
We investigate the accretion of a collisionless, relativistic kinetic gas by a rotating Kerr black hole, assuming that at infinity the state of the gas is described by a distribution function depending only on the energy of the particles. Neglecting the self-gravity of the gas, we show that relevant physical observables, including the particle current density and the accretion rates associated with the mass, the energy, and the angular momentum, can be expressed in the form of closed integrals that can be evaluated numerically or approximated analytically in the slow-rotation limit. The accretion rates are computed in this manner for both monoenergetic particles and the Maxwell-Jüttner distribution and compared with the corresponding results in the non-rotating case. We show that the angular momentum accretion rate decreases the absolute value of the black hole spin parameter. It is also found that the rotation of the black hole has a small but non-vanishing effect on the mass and the energy accretion rates, which is remarkably well described by an analytic calculation in the slow-rotation approximation to cubic order in the rotation parameter. The effects of rotation on the morphology of the accretion flow are also analyzed.
@@ I Introduction
At the same time, non-spherical flows in the Schwarzschild or Kerr spacetime were investigated within the same framework in [18, 19, 20, 21, 22] (axially symmetric disks), or [23, 24, 25, 26] (accretion onto moving black holes). A Bondi-type model confined to the equatorial plane of the Kerr spacetime was derived in [27] and [28]. Analytic kinetic models describing the relaxation (due to phase-space mixing) of macroscopic variables to a stationary state around a Kerr black hole have been described in [29, 30]. For self-gravitating shells or disks made of Vlasov matter around black holes, see [31, 32, 33, 34].
@@
We show that the particle number and energy accretion rates decrease with the increasing black hole spin parameter. This stays in agreement with previous results, in particular the model of accretion onto Reissner-Nordström black holes [11, 12] and the planar model discussed in [27]. This is also a general feature present in hydrodynamical models, e.g., in [5]. In contrast to the results of [27], both accretion rates decrease with increasing asymptotic temperature of the gas. This fact emphasizes the difference between planar and non-planar models, mostly due to their asymptotic properties. Further, we show that the angular momentum accretion rate is roughly proportional to the spin parameter, slowing the black hole rotation down. This stays in agreement with the results of [27], however it is not obvious—the hydrodynamical model of [1] is characterized by a vanishing angular momentum accretion rate.
Our solution is exact in the sense that the relevant observable quantities (such as the components of the particle current density or the components of the energy-momentum-stress tensor) can be expressed in the form of closed integrals. These integrals are then evaluated numerically. As an example, we compute the components of the particle current density, illustrating the morphology of the flow in the vicinity of the black hole.
@@ II Preliminaries
$\hat{L}_{z}:=L_{z}-aE,\qquad\beta:=\frac{\hat{L}_{z}}{L},$ (8)
since this leads to a simple form of the effective potential describing the radial motion [39, 4].
@@ III.2 New parametrization
    In the Schwarzschild limit $\alpha=0$, $Q_{c}=L_{c}(E)$, and $Q_{\mathrm{max}}=L_{\mathrm{max}}(r,E)$ are independent of $\alpha$, $\sin\chi$, and $\vartheta$ and reduce to the corresponding expressions in [9].
@@ III.5 Asymptotic conditions
$F_{\infty}=Am\delta(E-E_{0})=A\delta(\varepsilon-\varepsilon_{0}),$ (61)
where $A$ and $E_{0}=m\varepsilon_{0}$ are constants. In this case, the gas comprises of same-energy particles. Our second choice would be to set
@@ IV Accretion rates
The energy and angular momentum accretion rates can be obtained in the same fashion by inserting the factors $p_{t}=E$ and $p_{\varphi}=Q\sin\vartheta\sin\chi+a\sin^{2}\vartheta E$ under the integral (75), which yields
@@
$\mathcal{\dot{J}}=-2\pi\int\limits_{m}^{\infty}dEF_{\infty}(E)\int\limits_{0}^{\pi}d\vartheta\sin^{2}\vartheta\int\limits_{0}^{2\pi}d\chi\left[\frac{1}{3}\sin\chi Q_{c}^{3}+\frac{a}{2}\sin\vartheta EQ_{c}^{2}\right].$ (78)
@@
$\frac{\dot{\mathcal{N}}}{n_{\infty}}$ $=$ $\frac{\dot{\mathcal{E}}}{\varepsilon_{\infty}}=-\frac{M^{2}}{4\varepsilon_{0}\sqrt{\varepsilon_{0}^{2}-1}}\int\limits_{0}^{\pi}d\vartheta\sin\vartheta\int\limits_{0}^{2\pi}d\chi q_{c}(\alpha,\vartheta,\sin\chi,\varepsilon_{0})^{2},$ (80a)
$\frac{\dot{\mathcal{J}}}{\varepsilon_{\infty}}$ $=$ $-\frac{M^{3}}{2\varepsilon_{0}^{2}\sqrt{\varepsilon_{0}^{2}-1}}\int\limits_{0}^{\pi}d\vartheta\sin^{2}\vartheta\int_{0}^{2\pi}d\chi\left[\frac{1}{3}\sin\chi q_{c}(\alpha,\vartheta,\sin\chi,\varepsilon_{0})^{3}+\frac{\alpha}{2}\sin\vartheta\varepsilon_{0}q_{c}(\alpha,\vartheta,\sin\chi,\varepsilon_{0})^{2}\right],$ (80b)
@@
$\frac{\dot{\mathcal{J}}}{\varepsilon_{\infty}}$ $=$ $-\frac{M^{3}z}{2\left[K_{1}(z)+3K_{2}(z)/z\right]}$ (81c)
$\times\int\limits_{1}^{\infty}d\varepsilon\exp(-z\varepsilon)\int\limits_{0}^{\pi}d\vartheta\sin^{2}\vartheta\int\limits_{0}^{2\pi}d\chi\left[\frac{1}{3}\sin\chi q_{c}(\alpha,\vartheta,\sin\chi,\varepsilon)^{3}+\frac{\alpha}{2}\sin\vartheta\varepsilon q_{c}(\alpha,\vartheta,\sin\chi,\varepsilon)^{2}\right],$
@@
$\lim\limits_{z\to\infty}\frac{1}{\sqrt{2\pi z}}\frac{\dot{\mathcal{N}}}{n_{\infty}}$ $=$ $\lim\limits_{z\to\infty}\frac{1}{\sqrt{2\pi z}}\frac{\dot{\mathcal{E}}}{\varepsilon_{\infty}}=-\frac{M^{2}}{4\pi}\int\limits_{0}^{\pi}d\vartheta\sin\vartheta\int\limits_{0}^{2\pi}d\chi q_{c}(\alpha,\vartheta,\sin\chi,1)^{2},$ (82a)
$\lim\limits_{z\to\infty}\frac{1}{\sqrt{2\pi z}}\frac{\dot{\mathcal{J}}}{\varepsilon_{\infty}}$ $=$ $-\frac{M^{3}}{2\pi}\int\limits_{0}^{\pi}d\vartheta\sin^{2}\vartheta\int\limits_{0}^{2\pi}d\chi\left[\frac{1}{3}\sin\chi q_{c}(\alpha,\vartheta,\sin\chi,1)^{3}+\frac{\alpha}{2}\sin\vartheta q_{c}(\alpha,\vartheta,\sin\chi,1)^{2}\right].$ (82b)
In the high-temperature limit one uses the substitution $\varepsilon\mapsto\eta:=z\varepsilon$ and $K_{\nu}(z)\sim\frac{1}{2}\Gamma(\nu)(z/2)^{-\nu}$ in Eq. (81) and takes the limit $z\to 0$ to obtain
$\lim\limits_{z\to 0}\frac{\dot{\mathcal{N}}}{n_{\infty}}$ $=$ $\lim\limits_{z\to 0}\frac{\dot{\mathcal{E}}}{\varepsilon_{\infty}}=-\frac{M^{2}}{4}\int\limits_{0}^{\pi}d\vartheta\sin\vartheta\int\limits_{0}^{2\pi}d\chi\zeta_{c}(\alpha,\vartheta,\sin\chi)^{2},$ (83a)
$\lim\limits_{z\to 0}\frac{\dot{\mathcal{J}}}{\varepsilon_{\infty}}$ $=$ $-\frac{M^{3}}{2}\int\limits_{0}^{\pi}d\vartheta\sin^{2}\vartheta\int\limits_{0}^{2\pi}d\chi\left[\frac{1}{3}\sin\chi\zeta_{c}(\alpha,\vartheta,\sin\chi)^{3}+\frac{\alpha}{2}\sin\vartheta\zeta_{c}(\alpha,\vartheta,\sin\chi)^{2}\right],$ (83b)
@@ V Slow-rotation approximation
Since $Q_{c}=Q_{c}(a,\vartheta,\chi,E)$ is only given in parametric form (see Eq. (128)), the integrals with respect to $\chi$ present in Eqs. (45) cannot be computed analytically and one has to resort to numerical methods. Although we shall compute these integrals numerically in the upcoming section, here we circumvent this problem by considering a slowly rotating Kerr black hole and performing an expansion in the rotation parameter, up to cubic order. The slow-rotation approximation is an important astrophysically motivated case [50], and it provides an effective strategy to overcome difficulties arising in the Kerr background spacetime due to the frame-dragging effect [51]. In principle, the analytic formulas presented in this section can be extended to higher order of the black hole spin parameter, to derive relations valid for more rapidly rotating black holes. However, as we will see, the cubic approximation already leads to an accurate description of the accretion process.
Up to the third order in $\alpha$, the critical value $Q_{c}$ can be expressed explicitly in terms of $E$, $\chi$, $\vartheta$, and $\alpha$ in the following way. First, we expand the critical radius
$x_{c}=x_{c}^{(0)}+\alpha x_{c}^{(1)}+\alpha^{2}x_{c}^{(2)}+\alpha^{3}x_{c}^{(3)}+\mathcal{O}(\alpha^{4}),$ (85)
@@
which is a function of energy only. To find the correction terms $x_{c}^{(1)}$, $x_{c}^{(2)}$, and $x_{c}^{(3)}$, we substitute the ansatz (85) into the expression for $E_{c}$ given in Eq. (128) and then solve the equation $E_{c}=E$ order by order in $\alpha$. After some manipulations, this yields
@@
$x_{c}^{(3)}$ $=$ $\left(\frac{\sin\vartheta\sin\chi}{x_{c}^{(0)}-6}\right)^{3}\left[\frac{-7\left(x_{c}^{(0)}\right)^{2}+36x_{c}^{(0)}-12}{\sin^{2}\vartheta\sin^{2}\chi\left(x_{c}^{(0)}\right)^{3/2}}+\frac{-3\left(x_{c}^{(0)}\right)^{3}+23\left(x_{c}^{(0)}\right)^{2}-48x_{c}^{(0)}+12}{\sin^{2}\chi\left(x_{c}^{(0)}\right)^{3/2}}\right$. (87c)
$\left.+\frac{3\left(x_{c}^{(0)}\right)^{4}-43\left(x_{c}^{(0)}\right)^{3}+230\left(x_{c}^{(0)}\right)^{2}-516x_{c}^{(0)}+360}{\sqrt{x_{c}^{(0)}}\left(x_{c}^{(0)}-6\right)^{2}}\right],$
@@
Next, we substitute this function into the expression for $Q_{c}$ given in Eq. (128) and expand the result in powers of $\alpha$. In terms of the dimensionless quantities $q_{c}=Q_{c}/(Mm)$ this yields, up to the third order,
$q_{c}=q_{c}^{(0)}+\alpha q_{c}^{(1)}+\alpha^{2}q_{c}^{(2)}+\alpha^{3}q_{c}^{(3)}+\mathcal{O}(\alpha^{4}),$ (88)
@@
$q_{c}^{(2)}$ $=$ $\frac{q_{c}^{(0)}}{2x_{c}^{(0)}}\left[-\frac{\cos^{2}\vartheta+x_{c}^{(0)}\sin^{2}\vartheta}{x_{c}^{(0)}}+\frac{\left(x_{c}^{(0)}-5\right)\sin^{2}\vartheta\sin^{2}\chi}{x_{c}^{(0)}-6}\right],$ (89c)
$q_{c}^{(3)}$ $=$ $\frac{\sin^{3}\vartheta\sin^{3}\chi}{\left(x_{c}^{(0)}-6\right)\sqrt{x_{c}^{(0)}-3}}\left[\frac{x_{c}^{(0)}+2\cot^{2}\vartheta}{\sin^{2}\chi\left(x_{c}^{(0)}\right)^{3/2}}-\frac{\left(x_{c}^{(0)}\right)^{2}-10x_{c}^{(0)}+26}{\sqrt{x_{c}^{(0)}}\left(x_{c}^{(0)}-6\right)^{2}}\right].$ (89d)
Note that the zeroth order term again agrees with the value known from the Schwarzschild case, see Eq. (22). Together with Eq. (86), Eqs. (88) and (89) provide the desired expansion for $Q_{c}$ as a function of $E$, $\chi$, $\vartheta$, and $\alpha$.
@@
$\int_{0}^{\pi}d\vartheta\sin\vartheta\int_{0}^{2\pi}d\chi q_{c}^{2}=4\pi\frac{\left(x_{c}^{(0)}\right)^{2}}{x_{c}^{(0)}-3}\left[1-\frac{2\alpha^{2}}{\left(x_{c}^{(0)}\right)^{2}\left(6-x_{c}^{(0)}\right)}+\mathcal{O}(\alpha^{4})\right],$ (90)
@@
On the other hand, one can numerically perform the following integrations over the energy and compute the accretion rates for generic values of $z$ up to the third order in the rotation parameter $\alpha$. Using Eq. (90), one obtains
$\frac{\dot{\mathcal{N}}}{n_{\infty}}$ $=$ $-\frac{\pi zM^{2}}{K_{2}(z)}\int_{1}^{\infty}\frac{\left(x_{c}^{(0)}\right)^{2}}{x_{c}^{(0)}-3}\left[1-\frac{2\alpha^{2}}{\left(x_{c}^{(0)}\right)^{2}\left(6-x_{c}^{(0)}\right)}\right]e^{-z\varepsilon}d\varepsilon+\mathcal{O}(\alpha^{4}),$ (92a)
$\frac{\dot{\mathcal{E}}}{\varepsilon_{\infty}}$ $=$ $-\frac{\pi z^{2}M^{2}}{zK_{1}(z)+3K_{2}(z)}\int_{1}^{\infty}\frac{\left(x_{c}^{(0)}\right)^{2}}{x_{c}^{(0)}-3}\left[1-\frac{2\alpha^{2}}{\left(x_{c}^{(0)}\right)^{2}\left(6-x_{c}^{(0)}\right)}\right]\varepsilon e^{-z\varepsilon}d\varepsilon+\mathcal{O}(\alpha^{4}),$ (92b)
$\frac{\dot{\mathcal{J}}}{\varepsilon_{\infty}}$ $=$ $\frac{4\pi\alpha z^{2}M^{3}}{3[zK_{1}(z)+3K_{2}(z)]}\int_{1}^{\infty}\left(\frac{x_{c}^{(0)}}{x_{c}^{(0)}-3}\right)^{\frac{3}{2}}\left\{1+\frac{3\alpha^{2}\left(x_{c}^{(0)}-3\right)\left[\left(x_{c}^{(0)}\right)^{2}-24\right]}{5\left(x_{c}^{(0)}-6\right)^{3}\left(x_{c}^{(0)}\right)^{2}}\right\}e^{-z\varepsilon}d\varepsilon+\mathcal{O}(\alpha^{4}).$ (92c)
Therefore, within the slow-rotation approximation, the original triple integrals appearing in Eqs. (81) can be replaced with simpler formulas (92). Remarkably, for all numerical examples discussed in Sec. VI, the difference between the results of the slow-rotation approximation and exact numerical values computed according Eqs. (81) stays below $4\%$ for the mass accretion rate and below $5\%$ for the angular momentum accretion rate, even for rapidly rotating black holes with rotation parameters up to $\alpha\leq 0.99$.
@@
$\lim\limits_{z\rightarrow\infty}\frac{1}{\sqrt{2\pi z}}\frac{J_{t}^{\mathrm{(tot)}}}{n_{\infty}}$ $=$ $-\frac{1}{\pi}\sqrt{\frac{2}{x}}\left(1+\frac{\alpha^{2}}{8x}\right)+\mathcal{O}(\alpha^{3}),$ (94a)
@@
$\lim\limits_{z\rightarrow\infty}\frac{1}{\sqrt{2\pi z}}\frac{J_{\varphi}^{\mathrm{(tot)}}}{n_{\infty}}$ $=$ $0+\mathcal{O}(\alpha^{3}),$ (94c)
where up to second order in the rotation parameter, $J_{\varphi}^{\mathrm{(tot)}}$ vanishes in the low-temperature regime.
@@
$\lim\limits_{z\rightarrow\infty}\frac{1}{\sqrt{2\pi z}}\frac{n(x,\vartheta)}{n_{\infty}}=\frac{1}{4\pi x^{3}}\sqrt{\frac{x(x+2)+4}{2x}}\left[8x^{2}+\frac{\alpha^{2}(x+2)}{x(x+2)+4}\left(8\sin^{2}\vartheta+x^{2}-\frac{16}{x+2}\right)\right]+\mathcal{O}(\alpha^{3}).$ (95)
It is worth mentioning that Eq. (95) reduces to $\sqrt{3}/\pi$ at $x_{+}=2$ in the standard Schwarzschild limit $\alpha\to 0$ as follows from [9]. We would like to stress that other observables and physical quantities can be computed within the slow-rotation approximation in an analogous way; the limited examples provided in this work are intended to demonstrate this technique.